Sol Summary – January 17th

Crew 219 Sol Summary Report 17-01-2020

Sol: 12

Summary Title: Mission Complete.

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 12 started with gourmet coffee and a hearty breakfast of bacon and eggs. After breakfast D. Masaitis performed a walk-around inspection of the campus with Handbook in hand, to annotate any deficiencies that needed to be fixed before check-out, and passed the list to the crew. As the crew began touching up the modules and Hab, D. Masaitis sat down to truncate the rough draft of the crew’s Mission Summary from almost 7,000 words of eloquent academic prose. The crew had completed their cleaning tasks by lunchtime, and spent much of the afternoon working on the 2000 piece Mars puzzle. N. Hadland and H. Blackburn joined D. Masaitis in The Great Revision, and the Mission Summary was complete by 1630 hours.

Crew 219 ended Simulation at 1700 hours, and stepped out into the waning daylight to clean exterior areas, completing the work by sundown. We are proud of the work that we have accomplished here, and look forward to meeting Crew 220 tomorrow!

Look Ahead Plan: Meeting and Training Crew 220!

Anomalies in work: None

Weather: Chilly and sunny all day

Crew Physical Status: OUTSTANDING

EVA: NONE

Reports to be filed:

Sol Summary

Operations Report

GreenHab Report

Mission Summary (In lieu of Research Report)

Food Inventory

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested:

None

Sol Summary – January 16th

Crew 219 Sol Summary Report 16-01-2020

Sol: 11
Summary Title: If You Give a Martian a Cookie
Author’s name: Cynthia Montanez

Mission Status: Nominal

Sol Activity Summary: The morning started with crew 219 sipping on coffee and indulging in a light breakfast before they partook in their nominal tasks. To train C. Montanez and K. Crisman for their next field season, C. Montanez took the position as CO and K. Crisman took the position as XO for the day. Between 0908 to 1018, EVA #17 entered sim for sample return and came back to a delicious dish of pork sausage meatballs and spaghetti made by H. Blackburn.

Once everyone finished eating, each person started to work on finishing up their projects, post-mission report edits, and/or cleaning up parts of the Hab. Between 1402 to 1534, EVA #18 entered sim for the final sample return and came back to freshly made chocolate chip oatmeal cookies, brownies, and chai tea made by H. Blackburn. The aroma of the desserts made the Hab smell like grandma’s house, as the taste brought the crew back to Earth.

Finally, the day ended with H. Blackburn cooking coconut curry and crew positions being switched back. Overall, today was a great day on Mars.

Look Ahead Plan: Cleaning the Hab and its exterior facilities in preparation for the next crew

Weather: Chilly and sunny all-day

Crew Physical Status: Nominal

EVA:

EVA #17 – Sample Return W of Candor Chasma (0520000/4251500)
EVA #18 – Sample Return 1 km S of Kissing Camel (0518500/4248300)

Reports to be filed:

Sol Summary
EVA Reports #17 and #18
Operations Report
GreenHab Report
Research Report
Astronomy Report
Journalist’s Report
Photos (Including Photo of the Day)

Support Requested: None

Sol Summary – January 15th

Crew 219 Sol Summary Report 15-01-2020

Sol: 10

Summary Title: What day is it, camel?

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The crew filtered out casually for coffee and a
light breakfast, before EVA #15 moved downstairs to suit up. The team
moved out towards North Ridge, documenting their route up to Hab
Ridge, so that future teams can use the same route. They found a
multitude of viable samples and returned to the Hab in time to have a
quick lunch and help EVA #16 get out on time. While EVA #16 moved out
for sampling, those at the campus processed samples and started
drafting language for their respective sections of the end of mission
summary. H. Blackburn prepared cheddar biscuits and hot chocolate for
the crew to enjoy as EVA #16 returned, and crew members got to work
preparing their evening reports. H. Blackburn prepared stir-fry in the
wok for dinner, and after the Comms window, the crew looks forward to
turning command over to C. Montanez and K. Crisman for a Thursday of
on-the-job command experience so they will have a taste of what is to
come next year. D. Masaitis will spend the day assuming the role of
Health and Safety Officer, and N. Hadland will take over as GreenHab
Officer. Their Command-exchange will end after tomorrow’s Comms
window, but will better prepare them to understand the roles they will
be returning to fulfill next year.

Look Ahead Plan: Two sample return EVA’s, materials consolidation,
revisions of mission summary

Anomalies in work: None

Weather: Chilly and sunny all day

Crew Physical Status: Nominal

EVA:

EVA #15 – Sample Collection on Hab Ridge (0517700/4251600)

EVA #16 – Sample Collection around North Pinto Hills (0521000/4248800)

Reports to be filed:

Sol Summary

EVA Report #15 & 16

EVA Request #17 & 18

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested:

None

Sol Summary – January 14th


Crew 219 Sol Summary Report 14-01-2020

Sol: 9

Summary Title: Taco Tuesday!

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: There is no Tuesday better than Taco Tuesday.
The morning started with coffee and a light breakfast as the crew got
up and started performing their normal routines around the Hab. The
sun shone brightly and everyone’s mood was light and jovial, as crew
members busied themselves and joked amongst one another. H. Blackburn
prepared tuna tacos for lunch, to the great enjoyment of everyone
present, and then the EVA support team went downstairs to prepare
suits for EVA #14’s 1300 departure. Checklists were followed, and EVA
#14 exited the airlock a minute early. They returned an hour and a
half later with a bag full of samples, and the much of the crew spent
the afternoon in the science dome investigating and cataloging the new
scientific material, while K. Crisman collected his biometric data
from the EVA’s participants. The crew then sat down and wrote their
evening’s reports while H. Blackburn effortlessly whipped up a
delicious chicken pot pie...of which there were no leftovers. After
Comms window, the team will meet one more time to discuss tomorrow’s
plan before bedding down for another full day on Mars. The sun shines
on us again, and the planet is finally starting to peek some red
terrain out from under sheets of snow!

Look Ahead Plan: Two EVAs, sample analysis and preparation for
shipment, rough drafts of the end of mission summary and maintenance
as necessary

Anomalies in work: None

Weather: cold and sunny all day

Crew Physical Status: Nominal

EVA:

EVA #14 – Geologic Sample Collection south of Robert’s Rock Garden
(0518500/4248300)

Reports to be filed:

Sol Summary

EVA Report #14

EVA Request #15 & #16

Operations Report

GreenHab Report

Research Report

Astronomy Report

Journalist's Report

Photos (Including Photo of the Day)

Support Requested: None

Sol Summary – January 13th

  

 Crew 219 Sol Summary Report 13-01-2020

Sol: 8

Summary Title: Oh, Mondays...

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The crew was woken abruptly around 0600 to a
generator failure and low power, and immediately responded according
to guidance from Outpost. We then boiled water on the stove for
coffee, and got started with the days work. Some worked on science
while others transplanted plants in the GreenHab. H. Blackburn made a
lunch of chili and rice before EVA #13 pushed out towards Beige Moon,
and those left behind continued their tasks. EVA #13 collected their
samples, but suffered an H&S-related incident (see HSO report) forcing
their return by 1508. K. Crisman resolved the situation swiftly,
decisively, and admirably. With the whole crew back at nominal status,
A. Elnajdi and R. Raphael prepared an incredible dinner for the crew
while everyone finished and proofread reports for this evening’s Comms
window.

Look Ahead Plan: One EVA, continued sample analysis, and maintenance
as necessary

Anomalies in work: Generator Failure

Weather: Early morning fog, cold and sunny all day

Crew Physical Status: Nominal at present (See HSO Report)

EVA:

EVA #13 – Geologic Sample Collection around Beige Moon (0516000/4254300)

Reports to be filed:

Sol Summary

EVA Report #13

EVA Request #14

HSO Report

Operations Report

GreenHab Report

Research Report

Journalist's Report

Photos (Including Photo of the Day)

Support Requested:

-Quart-sized ziploc bags

-Medical tape

-Refer to Operations Report for maintenance-related request 

Sol Summary – January 12th

Crew 219 Sol Summary Report 12-01-2020

Sol: 7

Summary Title: Bacon, Burgers, Hawaiian Shirts and Mistaken Aces

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Despite today being a DONSA (see yesterday’s Sol
Summary for definition), work did actually get done today. After a
breakfast of bacon and eggs, some crew members swept snow out of the
tunnels, while others swept and mopped mud from the Hab’s first floor.
A. Elnajdi performed some sample analysis in the Science Dome and R.
Raphael performed astronomic data reduction while H. Blackburn and C.
Montanez made bacon cheeseburgers and fresh salad with GreenHab greens
for lunch. Afterwards N. Hadland and R. Raphael simulated a dust storm
for our dust mitigation experiments, and others engaged in a lively
game of Rummy 500. C. Montanez took a victory, before the crew sat
down to finish reports and sip the world’s best chai tea, courtesy of
A. Elnajdi. The evening will hold a crew meeting to prepare for
tomorrow’s work, and then it’s back to business in the morning.

Look Ahead Plan: One EVA, continued sample analysis, and maintenance
if necessary

Anomalies in work: None

Weather: Cold and overcast most of the day, with some sunshine in the
late afternoon

Crew Physical Status: Nominal

EVA:

None

Reports to be filed:

Sol Summary

EVA Request #13

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Astronomy Photos (ATTN: Dr. Zubrin)

Photos (Including Photo of the Day)

Support Requested:

– Refer to Operations Report

Sol Summary – January 10th

Crew 219 Sol Summary Report 10-01-2020

Sol: 5

Summary Title: Frozen Fog & Tang-glazed Cinnamon Rolls

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The morning broke with the entire valley
shrouded in thick fog, dropping visibility next to nothing. By the
time EVA #9 was intended to depart, the morning sunshine had burned
off much of the fog, but ground conditions did not appear to support
safe maneuver conditions, so we decided to cancel it. The crew instead
busied themselves with science and maintenance, while H. Blackburn
prepared a wonderful lunch of pork sausage chili and fresh-baked
cornbread. The team for EVA #10 ate lunch and moved out on schedule,
achieving both their maneuver plan and sampling schedule. On the way
back, they recovered R. Raphael’s experimental mirror emplacement for
analysis. Upon their return they found another surprise, courtesy of
H. Blackburn – steaming hot chocolate and Tang-glazed cinnamon rolls –
a real treat! With full bellies, some crew members took a nap, while
others refined their reports for the evening Comms window. This
evening will consist of relaxed socialization while crew members
recoup for another Sol of hard work tomorrow.

Look Ahead Plan: We anticipate two EVA’s, and will continue ongoing
scientific analysis and maintenance operations.

Anomalies in work: None

Weather: Cold and foggy in the early morning, followed by clear,
sunny, and frozen conditions the rest of the day

Crew Physical Status: Nominal

EVA:

EVA #9 – Cancelled at Commander’s Discretion

EVA #8 – Geologic Sample Collection South of Mid Ridge Planitia
(0516800/4251600)

Reports to be filed:

Sol Summary

EVA Report 10

EVA Requests 11 & 12

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested: Refer to Operations Report

Sol Summary – January 8th

  

Crew 219 Sol Summary Report 08-01-2020

Sol: 2

Summary Title: Pancakes, Fresh Bread, Sunshine, and Snow Flurries

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 3 began with N. Hadland and A. Perez preparing a delicious pancake breakfast for the crew, followed by the departure and early return of EVA #5. Afterwards, the crew settled into science and maintenance tasks, punctuated by periods of novel reading. By 1230 hours, all hands went below to help EVA#6 prepare for departure. While they were out H. Blackburn baked a couple loaves of fresh bread, which would be summarily consumed after the EVA’s return. The crew spent their time before Comms troubleshooting radios and thawing water pipes (again). In great news, Crew 219 now has well established troubleshooting procedures to keep water flowing from the static tank to the loft tank.

Look Ahead Plan: We look forward to two more sampling EVA’s and will continue maintenance and science work around the campus.

Anomalies in work: None

Weather: Cold and overcast all day, with temperatures below freezing. There was a brief moment of sunshine towards sundown, which was quickly replaced with cloud cover and light snow flurries.

Crew Physical Status: Nominal

EVA:

EVA #5 – Geologic Sample Collection SW of Hab Ridge (0517500/4250100)

EVA #6 – Geologic Sample Collection around Phobos Peak (0519500/4250100)

Reports to be filed:

Sol Summary

EVA Reports 5 & 6

EVA Requests 7 & 8

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested:

Outpost is tracking current Support Request

Sol Summary – January 07

Crew 219 Sol Summary Report 07-01-2020

Sol: 2

Summary Title: Research, Research, Research

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 2 marks our full plunge into scientific research. With our mission plan published, we spent the day starting and continuing experiments, and conducted two more EVA’s for sample collection. Other features of the day include fresh bread and cobbler by the Lead Science Officer, and whispers of a showing of The Martian after the Comms window.

Look Ahead Plan: Tomorrow holds plans for two more proposed sample collection EVA’s, maintenance on a couple EVA suits, continued sample processing in the science dome, and continuation of our other experiments.

Anomalies in work: None

Weather: Cold and overcast all day, with temperatures below freezing.

Crew Physical Status: Nominal

EVA:

EVA #3 – Geologic Sample Collection near North Ridge (0518600/4250900)

EVA #4 – Geologic Sample Collection around Kissing Camels (0518100/4249300)

Reports to be filed:

Sol Summary
EVA Reports 1 & 2
EVA Requests 3 & 4
Operations Report
GreenHab Report
Research Report
Journalist’s Report
Photos (Including Photo of the Day)

Support Requested:

(Carried over from previous day)

Sundries: Paper towels
Medical: C-Spine needs to be replaces (broken/misshapen)
12x Medium Gauze Pads
Anitbiotic Ointment

Sol Summary – January 06th

  

Crew 219 Sol Summary Report 06-01-2020

Sol: 1

Summary Title: "I have 5 high-def pictures of a toilet. Awesome!"

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 1 began with inclement weather and EVA delays at the Commander's discretion. Instead, Crew 219 spent the morning polishing our mission plan, conducting maintenance tasks around campus, reorganizing and consolidating sundries, and preparing materials for scientific work. After a lunchtime reevaluation of the weather, both EVAs were cleared to go, and both were conducted to standard. The late afternoon consisted of sample processing in the science dome, herb harvesting in the GreenHab, report writing in the hab, and the successful execution of a toilet procedure provided by the Director.

Look Ahead Plan: Tomorrow holds plans for two proposed sample collection EVA's, maintenance on a couple EVA suits, and continued sample processing in the science dome.

Anomalies in work: None

Weather: Cold and foggy with intermittent light flurries in the early morning, to cold and overcast in the afternoon, with temperatures still holding just above freezing.

Crew Physical Status: 1 minor infury (Non-EVA) annotated in today's HSO report

EVA: EVA #1 - Terrain Familiarization and Sample Collection at Marble Ritual (0518700/4250700)
EVA #2 - Terrain Familiarization and Sample Collection at Marble Ritual (0518700/4250700)

Reports to be filed:

Sol Summary
EVA Reports 1 & 2
EVA Requests 3 & 4
Operations Report
HSO Report
GreenHab Report
Research Report
Journalist's Report
Mission Plan
Updated HSO Pre-Mission Checklist
Photos (Including Photo of the Day)

Support Requested:

Sundries: Paper towels
Medical: C-Spine needs to be replaced (broken/misshapen)
12x Medium Gauze Pads
Anitbiotic Ointment